OpenStudy (anonymous):

Given ΔABC, m∠A = 50°, m∠B = 60°, and a = 7. Find b.

OpenStudy (anonymous):

A) 5.4 B) 6.2 C) 7.9 D) 9.0

rishavraj (rishavraj):

apply sine rule

rishavraj (rishavraj):

\[\frac{ a }{ \sin A } = \frac{ b }{ \sin B } = \frac{ c }{ \sin C }\]

OpenStudy (anonymous):

and how do I do that?

rishavraj (rishavraj):

see as \[\frac{ a }{ \sin A } = \frac{ b }{ \sin B }\] u got the values of a, <A and <B so u can find the value of sinA and sinB and plug in the equation and get b

OpenStudy (anonymous):

C) 7.9?

rishavraj (rishavraj):

good work :))

OpenStudy (anonymous):

can you help me with one more?

rishavraj (rishavraj):

ooops....i gtg get ready for college .,...post it though :)

rishavraj (rishavraj):

apply cosine rule